OpenStudy (anonymous):

(sorry i need a lot of help lol) (see pic(its gonna take me a min to get it up there))

OpenStudy (anonymous):

\[\sqrt{12}-3\sqrt{3}\] 12 perfect square is 4 square root it 2 so that would be \[4\sqrt{3}-3\sqrt{3}\] i thought it would be \[1\sqrt{3}\] but that's not a choice?

OpenStudy (johnweldon1993):

You almost had it ...but remember \[\sqrt{12} = \sqrt{4}\sqrt{3} = 2\sqrt{3}\] Remember you have to put it like that first...

OpenStudy (anonymous):

o okay then it would be \[2\sqrt{3}\] right?

OpenStudy (johnweldon1993):

Right...for the √12 part yeah.... so now we have \[\large 2\sqrt{3} - 3\sqrt{3} = ?\]

OpenStudy (anonymous):

\[4\sqrt{3}\] \[-\sqrt{3}\] \[3\sqrt{9}\] these are the choices @johnweldon1993

OpenStudy (johnweldon1993):

Mmhmm I see it :) lol \[2\sqrt{3} - 3\sqrt{3} = -1\sqrt{3}\] Is that -1 really necessary there? Nope not at all..it is just \[\large -\sqrt{3}\]

OpenStudy (anonymous):

ooops i made a simple mistake, uhh i make those often but when i see them i understand what i did lol thank you! (AGAIN)

OpenStudy (johnweldon1993):

haha don't worry about it :) Good work :)